OpenStudy (vera_ewing):

At equilibrium, the concentration of calcium ions in a calcium chromate (CaCrO4) solution is 1 x 10^–4 mol/L at 25°C. In the same solution, the concentration of chromate ions is also 1 x 10^–4 mol/L. What is the Ksp of calcium chromate? Show your work.

OpenStudy (vera_ewing):

@aaronq Please help!

OpenStudy (welshfella):

sorry I've forgotten this stuff

OpenStudy (aaronq):

Write the dissociation, then it's corresponding equilibrium equation (which is the Ksp equation). Plug in your values and solve

OpenStudy (vera_ewing):

What's the dissociation and corresponding equilibrium equation?

OpenStudy (aaronq):

the dissociation of CaCrO4 (in water)

OpenStudy (aaronq):

and by that i mean a chemical equation

OpenStudy (vera_ewing):

Okay, so what does the chemical equation look like? Can you set it up and then I solve?

OpenStudy (aaronq):

nah, you have to try it

OpenStudy (vera_ewing):

I don't know how :( Can you maybe explain it?

OpenStudy (aaronq):

CaCrO4 is an ionic compound, right? what ions make up CaCrO4?

OpenStudy (vera_ewing):

Not sure. It's covalent and ionic though, right?

OpenStudy (aaronq):

it's ionic only. You should be able to identify the ions here, if you can't look the compound up on wikipedia and read a little about it

OpenStudy (welshfella):

what i do remember is that Ca CrO4 disassociates in water to [Ca++] and [Cr2O4 --]

OpenStudy (welshfella):

and sp means solubility product so i guess Ksp is the product of Ca++ ions and CrO4-- ions

OpenStudy (vera_ewing):

Yep, I figured it out :) thanks

OpenStudy (welshfella):

so that 10-4 8 10^-4 = 10^-8

OpenStudy (welshfella):

oh Ok wd

OpenStudy (photon336):

CaCrO4 ----> Ca^2+ + CrO4^2- Ksp = [Ca^2+][CrO4^2-] Ksp = [x][x] Ksp = concentration of ions dissolved, like any K it's an equilibrium value, and it is temperature dependent. I'm assuming that the equilibrium is already reached and that the solution is saturated. B.c from my understanding that's what Ksp is. (Qsp = Ksp) Their concentrations are equal for both ions in the problem. So I think it's (1x10^-4)(1x10^-4) = Ksp. 10^-8